Which of the Following is a factored form of tthe epression 7x^2+10x-8
A. (x-1) (7x+8)
B. (x-4) (7x+2)
C. (x-8) (7x-1)
D. (x+2) (7x-4)
E. (x+4) (7x-2)

Found 2 solutions by fulio1, MathTherapy:
Answer by fulio1(4)

The answer is D
(x+2)(7x-4)=
x*7x= 7x^2 and x*-4= -4x
+2*7x= +14x and +2*-4 = -8
7x^2-4x+14x-8 ---combine like terms
7x^2+10x-8

Answer by MathTherapy(10552)

D. (x+2) (7x-4)
7x^2 + 14x - 4x - 8
7x(x + 2)- 4(x + 2)
(7x - 4)(x + 2)